1. A reaction vessel contains an equilibrium mixture of SO2, O2, and SO3. The reaction proceeds such that: 2SO2(g) + O2(g) ⇌ 2SO3(g)
The partial pressures at equilibrium are: PSO2=0.001111 atm PO2=0.002728 atm PSO3=0.0166 atm. Calculate Kp for the reaction.
2. A reaction vessel contains an equilibrium mixture of SO2, O2, and SO3. The reaction proceeds such that: 2SO2(g) + O2(g) ⇌ 2SO3(g)
The partial pressures at equilibrium are: PSO2=0.001714 atm PO2=0.001919 PSO3=0.0166 atm
3. A reaction vessel contains an equilibrium mixture of SO2, O2, and SO3. The reaction proceeds such that: 2SO2(g) + O2(g) ⇌ 2SO3(g)
The partial pressures at equilibrium are: PSO2=0.002318 atm PO2=0.001515 atm PSO3=0.0166 atm
1
Expert's answer
2020-04-02T10:15:39-0400

2SO2(g) + O2(g) = 2SO3(g)

Kp = "{\frac {P(SO_3)^2} {P(SO_2)^2P(O_2)}}"

1)P(SO2) = 0.001111 atm

P(O2) = 0.002728 atm

P(SO3) =0.0166 atm

Kp = "{\frac {P(SO_3)^2} {P(SO_2)^2P(O_2)}}={\frac {(0.0166)^2} {(0.001111)^2(0.002728)}}=" 81835.9 atm-1

2)P(SO2) = 0.001714 atm

P(O2) = 0.001919 atm

P(SO3) =0.0166 atm

Kp ="{\frac {P(SO_3)^2} {P(SO_2)^2P(O_2)}}={\frac {(0.0166)^2} {(0.001714)^2(0.001919)}}=" 48878.7 atm-1

3)P(SO2) = 0.002318 atm

P(O2) = 0.001515 atm

P(SO3) =0.0166 atm

Kp ="{\frac {P(SO_3)^2} {P(SO_2)^2P(O_2)}}={\frac {(0.0166)^2} {(0.002318)^2(0.001515)}}=" 33851.4atm-1